4.What pressure (Pa) will carbon monoxide ( II ) create with a mass of 140 g at a temperature of 40 ° C in a container with a volume of 200 liters?


1
Expert's answer
2021-06-11T22:47:54-0400

There are two scenarios for consideration in this question.

CASE 1: THE IDEAL GAS SCENARIO

for an ideal gas, it obeys the equation PV=nRT. we therefore make P the subject of this formula.

"P= \frac{n \times R \times T}{V}" where P is the pressure, V is the volume, T is the absolute temperature and R is the universal gas constant. The choice for R should be in conformity with the variables in use. for this case R=8314.46 L.pa.K-1mol-1 is the most convenient.

"\frac{((273 + 40) × (140 ÷ 44) × 8314.46)}{200} =41402.2315pa"

CASE 2: THE REAL GAS SCENARIO

For a real Gas, "P={nRT \over V-nb}- {an^2 \over V^2}" a and b for CO is 1.505 and 0.03985 respectively.

"{((3.182 × 8314.46 × 313) \over (200 - (3.182 × 0.03985)}-{1.505 \times 3.182^2 \over 200^2}=41430.8646886pa"

tis volume is slightly different from that of the ideal pressure due to correction factors for infinite volume and zero molecular attraction